OpenStudy (anonymous):

9x^3-43x^2-10x solve and show work please. Im stuck and need help please

OpenStudy (anonymous):

hold on let me try to figure this out.. I think I know how... give me a minute

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

youre working on solving polynomials right?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Are you solving\[9x^3-43x^2-10x=0\]? Do you see anything you can factor out there?

OpenStudy (anonymous):

umm yea the 9 and x

OpenStudy (whpalmer4):

well, \(x\) anyhow. \[x(9x^2-43x-10)=0\]That means that \(x=0\) is one solution, right?

OpenStudy (whpalmer4):

That leaves us with \[9x^2-43x-10=0\]You can factor that, or use the quadratic formula. Either way you should get two more solutions.

OpenStudy (anonymous):

i would rather factor that, right

OpenStudy (whpalmer4):

what do you get when you factor it? you must be a much better factoring fiend than I am :-)